Equivalent Units of Production: Average Cost Method

The following information concerns production in the Finishing Department for May. The Finishing Department uses the average cost method.

ACCOUNT Work in Process—Finishing Department ACCOUNT NO.
Date Item Debit Credit Balance
Debit Credit
May 1 Bal., 15,500 units, 70% completed 62,000
31 Direct materials, 76,000 units 185,440 247,440
31 Direct labor 115,000 362,440
31 Factory overhead 116,800 479,240
31 Goods transferred, 80,600 units 461,032 18,208
31 Bal., ? units, 35% completed 18,208

a. Determine the number of units in work in process inventory at the end of the month.
units

b. Determine the number of whole units to be accounted for and to be assigned costs and the equivalent units of production for May.

Whole units units
Equivalent units of production units

Cost of Production Report

The debits to Work in Process—Roasting Department for Morning Brew Coffee Company for August, together with information concerning production, are as follows:

Work in process, August 1, 900 pounds, 50% completed $3,465*
*Direct materials (900 X $3.2) $2,880
Conversion (900 X 50% X $1.3) 585
$3,465
Coffee beans added during August, 28,000 pounds 88,200
Conversion costs during August 38,458
Work in process, August 31, 1,400 pounds, 30% completed ?
Goods finished during August, 27,500 pounds ?

All direct materials are placed in process at the beginning of production.

a. Prepare a cost of production report, presenting the following computations:

  1. Direct materials and conversion equivalent units of production for August.
  2. Direct materials and conversion costs per equivalent unit for August.
  3. Cost of goods finished during August.
  4. Cost of work in process at August 31.

If an amount is zero, enter in "0". For the cost per equivalent unit, round your answer to two decimal places.

Morning Brew Coffee Company
Cost of Production Report-Roasting Department
For the Month Ended August 31
Unit Information
Units charged to production:
Inventory in process, August 1
Received from materials storeroom
Total units accounted for by the Roasting Department
Units to be assigned costs:
Equivalent Units
Whole Units Direct Materials (1) Conversion (1)
Inventory in process, August 1
Started and completed in August
Transferred to finished goods in August
Inventory in process, August 31
Total units to be assigned costs
Cost Information
Cost per equivalent unit:
Direct Materials Conversion
Total costs for August in Roasting Department $ $
Total equivalent units
Cost per equivalent unit (2) $ $
Costs assigned to production:
Direct Materials Conversion Total
Inventory in process, August 1 $
Costs incurred in August
Total costs accounted for by the Roasting Department $
Costs allocated to completed and partially completed units:
Inventory in process, August 1 balance $
To complete inventory in process, August 1 $ $
Cost of completed August 1 work in process $
Started and completed in August
Transferred to finished goods in August (3) $
Inventory in process, August 31 (4)
Total costs assigned by the Roasting Department $

b. Compute and evaluate the change in cost per equivalent unit for direct materials and conversion from the previous month (July). If required, round your answers to the nearest cent.

Increase or Decrease Amount
Change in direct materials cost per equivalent unit $
Change in conversion cost per equivalent unit $
Answer #1
1
a
Number of units in work in process inventory at the end of the month = 15500+76000-80600= 10900
b
Whole units = 15500+76000= 91500 units
Equivalent units of production = 80600+(10900*35%)= 84415 units
2
Morning Brew Coffee Company
Cost of Production Report-Roasting Department
For the Month Ended August 31
Unit Information
Units charged to production:
Inventory in process, August 1 900
Received from materials storeroom 28000
Total units accounted for by the Roasting Department 28900
Units to be assigned costs:
Equivalent Units
Whole Units Direct Materials Conversion
Inventory in process, August 1 900 0 450
Started and completed in August 26600 26600 26600
Transferred to finished goods in August 27500 26600 27050
Inventory in process, August 31 1400 1400 420
Total units to be assigned costs 28900 28000 27470
Cost Information
Costs per equivalent unit:
Direct Materials Conversion
Total costs for August in Roasting Department 88200 38458
Total equivalent units 28000 27470
Cost per equivalent unit 3.15 1.40
Costs charged to production:
Direct Materials Conversion Total
Inventory in process, August 1 3465
Costs incurred in August 126658
Total costs accounted for by the Roasting Department 130123
Cost allocated to completed and partially completed units:
Inventory in process, August 1 balance 3465
To complete inventory in process, August 1 0 630 630
Cost of completed August 1 work in process 4095
Started and completed in August 83790 37240 121030
Transferred to finished goods in August 125125
Inventory in process, August 31 4410 588 4998
Total costs assigned by the Roasting Department 130123
b
Increase or Decrease Amount
Change in direct materials cost per equivalent unit Decrease 0.05 =3.2-3.15
Change in conversion cost per equivalent unit Increase 0.10 =1.4-1.3
Answer #2

 

Beginning work in process   inventory

6,000 items, 1/4 complete

Units placed in production

14,000 units

Units completed

17,000 units

Ending work in process inventory

3,000 items, 50% complete



Cost of beginning work in process

P36,000

Direct material costs, current

  42,000

Conversion costs, current

P68,000

 

Direct materials are placed into production at the beginning of the process and conversion costs are incurred evenly throughout the process.

 

Required:

Prepare a production cost worksheet using the FIFO method.


source: ACCOUNTING
answered by: csella.avila123@gmail.com